OpenStudy (anonymous):

A team in science class placed a chalk mark on the side of a wheel and rolled the wheel in a straight line until the chalk mark returned to the position. The team then measured the distance of the wheel had rolled and found it to be 10 cm. to the nearest tenth, what is the area of the wheel. answer choices: a. 31.8cm^2 b. 15.9 cm^2 c.8 cm^2 d. 7.9cm^2

OpenStudy (anonymous):

@Abbles @TheRealMeeeee @McDonaldsGirl<3

OpenStudy (abbles):

The circumference is 10 cm then. C = dpi

OpenStudy (anonymous):

Yeah, all you need to do is use the circumference formula lol

OpenStudy (anonymous):

C = 2 * pi * r, r = 10

OpenStudy (anonymous):

err wait C = 10

OpenStudy (anonymous):

You solve for r

OpenStudy (abbles):

10 = d3.14 10/3.14 = d d = 3.18

OpenStudy (anonymous):